HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 6 of 8  Not logged in ELOG logo
New entries since:Thu Jan 1 01:00:00 1970
Entry  Wed Apr 2 02:51:31 2025, Carlo, Zac, Chris, DAQ, Beam is back 
Entry  Wed Apr 2 01:46:52 2025, Chris, Carlo, Zac, General, Beam loss 
Entry  Wed Apr 2 00:59:43 2025, Zac, Carlo, Chris, Accelerator, Lost the beam 
Entry  Wed Apr 2 00:08:57 2025, Ruijiu Chen,Shahab, Analysis, the peak corresponding to 194Pb⁸²⁺ was mixed with signals from some unknown ions. photo_2025-04-02_00-12-38.jpgphoto_2025-04-02_00-12-43.jpg
Entry  Tue Apr 1 23:56:18 2025, Shahab, Carlo, Zac, Yue, Michael, DAQ, Change of RSA01 settings for new harmonic=212. 
Entry  Tue Apr 1 23:23:18 2025, C Griffin, General, PLEIADES removed 
Entry  Tue Apr 1 23:07:15 2025, Shahab, DAQ, short interruption of queqiao script for RSAs 
Entry  Tue Apr 1 22:53:59 2025, Esther, General, Lassie exports 2025-04-01_22-53-17-548_tcl1146.png2025-04-01_22-53-21-680_tcl1146.tdf
Entry  Tue Apr 1 22:05:24 2025, Ruijiu , Analysis, analysis result 202504012157.png202504012157.pdf202504012206.pngphoto_2025-04-01_22-43-41.jpg
Entry  Tue Apr 1 22:03:07 2025, Jeroen, Zac, Esther, Shahab, Wolfram, DAQ, Optimisation of settings 
Entry  Tue Apr 1 20:37:04 2025, Jeroen, DAQ, Change of settings for shorter acq. window 
    Reply  Tue Apr 1 21:05:20 2025, Jeroen, Zac, Esther, Shahab, Wolfram, DAQ, Change of settings for shorter acq. window 
Entry  Tue Apr 1 20:00:14 2025, Jeroen, Zac, Esther, Accelerator, Beam fell out 
    Reply  Tue Apr 1 20:02:19 2025, Jeroen, Zac, Esther, Accelerator, Beam fell out 
       Reply  Tue Apr 1 20:19:23 2025, Jeroen, Zac, Esther, Accelerator, Beam fell out 2025-04-01_20-17-10-651_tcl1064.png
Entry  Tue Apr 1 17:08:13 2025, JG, ID, MC, CG, Detectors, pleiades in position 
Entry  Tue Apr 1 16:47:32 2025, Ruijiu, Jan, Shahab, Analysis, The frequency of 194Pb82+ is changed to 408.600 MHz Screenshot_2025-04-01_16-46-56.pngphoto_2025-04-01_16-54-20.jpg
Entry  Tue Apr 1 14:54:24 2025, all, General, No SIS beam since 14:35 
    Reply  Tue Apr 1 16:44:32 2025, Jeroen, Esther, General, No SIS beam since 14:35 2025-04-01_16-47-13-677_tcl1146.png2025-04-01_16-47-10-870_tcl1146.tdf
Entry  Tue Apr 1 15:38:50 2025, ruijiu chen, Analysis, frequency drift Screenshot_2025-04-01_15-38-26_1.pngScreenshot_2025-04-01_15-41-47.png
Entry  Tue Apr 1 15:32:59 2025, Zac, DAQ, NTCAP Restarted 
Entry  Tue Apr 1 14:25:21 2025, ruijiu , Analysis, The frequency of 194Pb82+ has been changed from 408.606MHz to 408.602MHy during last 12 hours 12x
Entry  Tue Apr 1 14:23:01 2025, JG, General, Pleiades pocket position tests 
Entry  Tue Apr 1 14:17:07 2025, Ruijiu, Analysis, PID of 194Pb82+ 2025.04.01.14.13.07 Screenshot_2025-04-01_14-16-42.png
ELOG V3.1.5-fc6679b